OpenStudy (anonymous):

which relation is not a function? 1) y=2x+4 2) y=x^2-4x+3 3) x=3y-2 4) x= y^2+2x-3

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (anonymous):

its neither 1 or 3 i m stuck between 2 and 4

OpenStudy (anonymous):

lol

OpenStudy (skullpatrol):

In math we start with definitions, right?

OpenStudy (skullpatrol):

What is the definition of a function?

OpenStudy (anonymous):

for every x there is only 1 y

OpenStudy (skullpatrol):

Yes...

OpenStudy (skullpatrol):

look at the form of the equations given

OpenStudy (skullpatrol):

for your idea to work they should be written how?

OpenStudy (skullpatrol):

think about it

OpenStudy (skullpatrol):

1) y=2x+4 2) y=x^2-4x+3 3) x=3y-2 4) x= y^2+2x-3

OpenStudy (skullpatrol):

You said "for every x there is only 1 y"

OpenStudy (anonymous):

yea

OpenStudy (skullpatrol):

so that means what?

OpenStudy (anonymous):

should i inverse choice 3 and 4 or something

OpenStudy (anonymous):

i believe 4) is not a function

OpenStudy (skullpatrol):

good you are getting the idea,,,choice 3 is OK because you can easily write as y = ... right?

OpenStudy (anonymous):

yes

OpenStudy (skullpatrol):

@amrita000 come on you can do it on your own

OpenStudy (skullpatrol):

try

OpenStudy (anonymous):

4) is not a function

OpenStudy (nincompoop):

how s square root not a function?

OpenStudy (anonymous):

x= y^2+2x-3, y^2=-x+3, y=sqrt(-x+3), if x is bigger than 3 then it is an imaginary function

OpenStudy (anonymous):

This is not true @AmardeepSingh

OpenStudy (anonymous):

its |y|=sqrt(3-x) and that's the difference (the absoute value)

OpenStudy (skullpatrol):

The question should read which relation is not a function of x @amrita000

OpenStudy (skullpatrol):

$$\Huge y=f(x)$$

OpenStudy (anonymous):

alright thanks .. i got it... its choice 4

OpenStudy (skullpatrol):

Good work :-)

OpenStudy (anonymous):

Thanks everybodyy!!! ^_^

OpenStudy (nincompoop):

knowing the graph of the functions would help a great deal, because a function passes the vertical line test

OpenStudy (anonymous):

@nincompoop +1 But it also takes more time!

OpenStudy (nincompoop):

disagree, it is the easiest

OpenStudy (skullpatrol):

@amrita000 remember the definition

OpenStudy (skullpatrol):

$$\Huge y=f(x)$$

OpenStudy (anonymous):

okiee :D

OpenStudy (skullpatrol):

one x goes in and one y comes out

OpenStudy (anonymous):

@nincompoop so she has trouble finding which of the expressions is not a function and she would be able to graph the last expression? If she was able to graph it, she wouldn't need help to recognize it wasnt a function in my opinion

OpenStudy (anonymous):

Anyway good job!

OpenStudy (skullpatrol):

Graphs are great once the student understands the logical links the definition has to the graph.

OpenStudy (anonymous):

Exactly!